OpenStudy (acespeedfighter):

Which of the following is the solution to the equation 4/5n = 20? (3 points) n = 4 n = 12 n = 16 n = 25

OpenStudy (riff-raff):

n=25

MARC (marc_d):

Step 1: Multiply 5 on both sides

MARC (marc_d):

\[\frac{ 4n }{ 5 }=20\]

MARC (marc_d):

\[\frac{ 4n }{ 5 }\times5=20\times5\]

MARC (marc_d):

\[\frac{ 4n }{ 1\cancel{5} }\times1\cancel{5}=100\]

MARC (marc_d):

\[4n=100\]

MARC (marc_d):

Step 2: Divide both sides by 4

MARC (marc_d):

\[\frac{ 4n }{ 4 }=\frac{ 100 }{ 4 }\]

MARC (marc_d):

\[n=\frac{ 100 }{ 4 }\]

MARC (marc_d):

\[n=?\]

MARC (marc_d):

solve it! @AceSpeedFighter